# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v4

@@ 7.4 Foundation Models assisting for Reinforcement Learning
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function. FAC (Ye et al. (2023a)) proposes using knowledge from foundation models as policy prior knowledge to improve sampling action efficiency, as value prior knowledge to measure the values of states and as success-reward prior knowledge to provide final feedback on task success.
@@
Following Fig. 7 and Tab. 1, pre-trained foundation models can engage with Policy modules in various forms. However, no benchmark indicates which approach is the most effective at present. Regarding end-to-end RFMs, there are still many architectural considerations to make models more interpretable and reliably stable.
Table 1: Robotic-specific Foundation Models.This table organizes information about robotic-specific foundation models, specifically focusing on robotic manipulation tasks. In the ‘dataset’ section, only manipulation datasets are considered. Four collection methods are defined in ‘Collection Method’: Teleoperation (Human control a robot to perform tasks from a distance, such as the use of VR headsets with controllers), videos (Tasks performed by other robots or humans are recorded for the robot to mimic.), simulation (Tasks are collected in a simulation environment), and skill libraries (Pre-trained models are used to execute in new scene to enhance the dataset with the check module to recognize the task success). ‘Size’ refers to the parameters of the model. ”Input Modality” refers to the modality in which data is inputted into the model. ‘Architecture’ refers to the structure of the model. ‘Hardware’ refers to the devices used during the model training stage. ‘Output modality’ shares the same definition in Section 7. ’Frequency’ refers to the rate at which the model inference and control. ‘Benchmark’ refers to standard tests, focusing solely on manipulation, used to measure the performance.
Paper Robotics Dataset Collection Method Size Input Modality Architecture Hardware Output Modality Frequency Benchmark
BC-Z
(Jang et al. (2022))
BC-Z dataset
(Jang et al. (2022))
Human Demonstrations1
Human Videos
Robot Videos
Image
Language or Video
ResNet18 encoder +
FiLM layers +
a two-layer MLP
18 NVIDIA
V100 GPUs
Target Pose 10HZ
Own benchmark
(100 tasks in total)
MOO
(Stone et al. (2023))
RT-1 dataset
(Brohan et al. (2022))
+ Self Collection
Human Demonstrations  [^1] M
Image
Language
RT-1 + OWL-ViT - Delta Pose -
Own benchmark
(1472 real world valuations)
RT-1
(Brohan et al. (2022))
RT-1 dataset
(Brohan et al. (2022))
Teleoperation 35M
Image
Language
Conditioned EfficientNet +
TokenLearnr +
Decoder-only transformer
- Delta Pose 3HZ
Own benchmark
(over 700 instructions)
RT-2
((Brohan et al. (2023)))
RT-1 dataset
(Brohan et al. (2022))
Teleoperation 3/5/12/55B
Image
Language
PaLI-X
(Chen et al. (2022)),
PaLM-E
(Driess et al. (2023))
Multi-TPU
Cloud Service
Delta Pose 1-3HZ
Language Table
Lynch et al. (2023)
Own benchmark
RT-X
(Padalkar et al. (2023a))
Open X-Embodiment
(Padalkar et al. (2023a))
Human Demonstrations -
Image
Language
PaLI-X
(Chen et al. (2022)),
PaLM-E
(Driess et al. (2023))
Multi-TPU
Cloud Service
Delta Pose 3-10Hz
Own benchmark
(3600 evaluation trails)
RT-H
(Belkhale et al. (2024))
Diverse+Kitchen
(Belkhale et al. (2024))
Human Demonstrations -
Image
Language
PaLI-X
(Chen et al. (2022))
- Delta Pose -
Own benchmark
(Kitchen tasks)
LEO
(Huang et al. (2023a))
CLIPort
(Shridhar et al. (2021))
Human Demonstrations 1.3/7/13B
Image
3D Point Cloud
Language
Decoder-only transformer
4 NVIDIA
A100 GPUs
Target Pose 1Hz
Own benchmark
(3 tasks for manipulation)
RoboFlamingo
(Li et al. (2023d))
CALVIN Bench
(Mees et al. (2022))
Teleoperation 3/4/9B
Image
Language
OpenFlamingo
(Awadalla et al. (2023))
8 NVIDIA
Tesla A100 GPUs
Delta Pose -
CALVIN Bench
(Mees et al. (2022))
Robocat
(Bousmalis et al. (2023))
Meta-World
(Yu et al. (2020))
DeepMind Control Suite
(Tassa et al. (2018))
+Self Collection
Simulation
Skill Library
364M/1.18B
Image
Language
Proprioception
Decoder-only transformer -
Delta Pose
Hindsight Image
10-20HZ
RGB-Stacking Benchmark
(Lee et al. (2021))
Roboagent
(Bharadhwaj et al. (2023))
RoboSet(MT-ACT)
(Bahl et al. (2023))
Teleoperation
Skill Library
Image
Language
Proprioception
Transformer One 2080Ti GPU Delta Pose 5Hz
RL BENCH
(James et al. (2020))
VIMA
(Jiang et al. (2023))
VIMA dataset
(Jiang et al. (2023))
Simulation 200M
Image
Language
Decoder-only transformer
8 NVIDIA
V100 GPUs
Target pose -
VIMA-BENCH
(Jiang et al. (2023))
MIDAS
(Li et al. (2023c))
VIMA dataset
(Jiang et al. (2023))
Simulation 92M
Image
Language
Decoder-only transformer
8 NVIDIA
A100 GPUs
Target pose -
VIMA-BENCH
(Jiang et al. (2023))
PaLM-E
(Driess et al. (2023))
PaLM-E dataset
(Driess et al. (2023))
Skill Library 12/22/562B
Image
Neural 3D Representations
States of robot or objects
Decoder-only transformer
Multi-TPU
Cloud Service
Delta Pose
Text
1-5Hz
Language Table
(Lynch et al. (2023))
Saycan
(Ahn et al. (2022))
Gato
(Reed et al. (2022))
Meta-World
(Yu et al. (2020))
and other 15 datasets
Simulation 1.2B
Image
Language
Proprioception
Decoder-only transformer TPU v3 Delta Pose 20HZ
Meta-World
(Yu et al. (2020))
GR-1
(Wu et al. (2023a))
Ego4D dataset
(Grauman et al. (2022))
CALVIN Bench
(Mees et al. (2022))
+Self Collection
Human Videos
Teleoperation
195M
Image
Language
Proprioception
Decoder-only transformer -
Delta Pose
Image
CALVIN Bench
(Mees et al. (2022))
QT-Opt
(Kalashnikov et al. (2018))
QT-Opt dataset
(Kalashnikov et al. (2018))
Simulation -
Image
Gripper Status
Height
Reinforcement Learning
10 NVIDIA
P100 GPUs
Target Pose -
Own benchmark
(28 test objects)
RLS
(Herzog et al. (2023b))
RLS dataset
(Herzog et al. (2023b))
Skill Library
(Simulation + Real World)
Image
Proprioception
CNN TPUv3 Target Pose -
Own benchmark
(9 scenarios)
Q-Transformer
(Chebotar et al. (2023b))
Q-Transformer dataset
(Chebotar et al. (2023b))
Human Demonstrations
Skill Library
Image
Language
Q-Transformer network - Target Pose 3HZ
RT-1
(Brohan et al. (2022))
OpenVLA
(Kim et al. (2024))
Open X-Embodiment
(Padalkar et al. (2023a))
DROID
(Khazatsky et al. (2024))
Human Demonstrations1
Teleoperation
7B
Image
Language
Llama 2 7B
64 NVIDIA
A100 GPUs
Delta Pose 6HZ
BridgeData V2
(Walke et al. (2023))
RT-1, RT-2
(Brohan et al. (2022, 2023))
3D Diffuser Actor
(Ke et al. (2024))
RLBench
(James et al. (2020))
CALVIN
(Mees et al. (2022))
Simulation
Teleoperation
RGBD
Language
Diffusion Model
Target Pose 1.67HZ
RLBench
(James et al. (2020))
CALVIN
(Mees et al. (2022))
DNAct
(Yan et al. (2024))
RLBench
(James et al. (2020))
Simulation
3D Point Cloud
Language
Diffusion Model
2 NVIDIA
RTX 3090 GPUs
Target Pose -
RLBench
(James et al. (2020))
  • 1
    The articles mention the use of human demonstration but do not provide detailed descriptions of the specific methods employed.
    “-” donates no data is reported.
@@ 8 Manipulation Data Generation
To propel robots into the era of general manipulation, the acquisition of vast amounts of data is indispensable (Padalkar et al. (2023b)). Collecting real-world data requires a lot of human labor and expensive remote teleoperation equipment. There are currently two methods for data collection: the bottom-up approach and the top-down step-by-step approach. The bottom-up approach focuses on having the robot perform trajectories first. Then, it uses methods like crowd-sourcing to label the data. The top-down step-by-step approach involves decision-makers setting task labels. The robot then performs tasks according to these labels. RoboVQA (Sermanet et al. (2023)) shows that the bottom-up approach is more efficient in data collection compared to the top-down step-by-step approach. It uses the collected data to train VideoCoCa, which helps in handling long-horizon tasks with a hierarchical method. DIAL (Xiao et al. (2022a)) uses a fine-tuned CLIP to replace humans in labeling robot trajectories during bottom-up data collection. This transforms the robot manipulation dataset on the internet into the robot-language manipulation dataset. PAFF (Ge et al. (2023)) points out that incorrect robot trajectories can be linked to new tasks and uses fine-tuned CLIP to label the incorrect robot trajectories with appropriate task labels.
@@
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023c)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge. Regardless of whether it’s in a real or simulated environment, improving the efficiency of the existing dataset is essential. The mainstream approach is dataset augmentation.
@@ 8.1 Simulator
The current mainstream simulators (Zhou et al. (2023)) include PyBullet (Coumans and Bai (2016)), MuJoCo (Todorov et al. (2012)), CoppeliaSim (Rohmer et al. (2013)), NVIDIA Omniverse and Unity. Pybullet is easy to use and integrate, but its graphics are quite basic. It is not suitable for applications that require complex visual effects. Therefore, Pybullet is often used together with Blender (Shi et al. (2024)). Mujoco offers a high-precision physics engine. It is suitable for simulating articulated and deformable object manipulation. However, it has a high entry barrier for beginners. CoppeliaSim offers a wide range of ready-made environments, objects, and prototyping robotic systems for users. However, when dealing with many robots or complex scenes, CoppeliaSim may encounter performance issues. NVIDIA Omniverse provides real-time physics simulation and realistic rendering. However, it requires significant computational resources. NVIDIA Omniverse offers many interfaces. Users can use these to develop various applications. For example, Issac Gym is a platform for robot reinforcement learning, developed using Omniverse. Unity offers rich visual effects and a user-friendly interface. It allows for the creation of highly interactive applications. However, its physics engine is still not precise enough. The basic components of a simulator are the physics engine and the renderer. Improvements in these components can enhance the capability of sensors in simulations, such as optical tactile sensors (Chen et al. (2023d)). We also hope the simulators can add sound engine and other features. This would make the simulated world feel more real. Learning-based simulators also show great potential. For example, Sora (Brooks et al. (2024)) and UniSim (Yang et al. (2023b)) use vast amounts of data from the internet to simulate the visual effects of many different actions.
@@ 8.2 Scene and Demonstration Generation
The Real-to-Sim method directly refers to a digital twin. The Real-to-Sim method utilizes 3D-reconstruction technology (in Sec. 6) or inverse graphics (Chen et al. (2024c)) to create the real-world scene in a virtual environment (Torne et al. (2024)). But, 3D reconstruct scene is static environment where objects lack real-world physical properties, such as material, mass and friction coefficients, and are non-interactive (Torne et al. (2024)). The inverse graphics method, such as URDFormer (Chen et al. (2024c)), directly generates interactive simulation environment and articulated objects from input RGB image. Compared to 3D-reconstruction methods, it reduces human involvement and produces interactive simulation environment. However, it lacks physical plausibility and fails to address the mismatch between the generated object’s physical properties and the real world. Previous research on physical parameter identification of real-world objects can be categorized into estimation from interaction (Seker and Kroemer (2024); Bohg et al. (2017); Xu et al. (2019)), estimation from demonstrations (Torne et al. (2024)). PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to estimate physical object attributes, such as material, from observation.
@@
As for the application of foundation models in Real-to-Sim methods, GRS (Zook et al. (2024)) employs SAM2 for object segmentation from RGBD image and utilizes VLMs to describe and match objects with simulation-ready assets. This approach combines the strengths of 3D-reconstruction and inverse graphics methods. It ensures the credibility of 3D-reconstruction methods and allowing objects in the scene to interact. However, it is impossible for the assert dataset to fully cover objects in the real world. Constructing an interact assert dataset often requires manual design by the creator or human-assisted interactive object generation. ACDC (Dai et al. (2024)) defines a digital cousin concept. Unlike a digital twin, it does not directly replicate a real-world counterpart. However, it retains similar geometric and semantic features by using similar asserts when the assert dataset does not include real-world objects. As for object pose, depth cameras are commonly used, but they struggle to capture reflective surfaces accurately. This limits their use in the wild. To address this issue, ACDC uses Depth-Anything-v2 (Yang et al. (2024a)), a state-of-the-art monocular depth estimation model, to estimate the depth map.
@@
The automated generation methods can be categorized into rule-based and learning-based approaches. For instance, ProcTHOR (Deitke et al. (2022)) introduces a procedural generation pipeline for interactive scenes using rule-based constraints and statistical priors. However, the generated scenes often rely on pre-defined priors, resulting in unrealistic outcomes that hinder agent learning (Khanna et al. (2024)). To address this, PHYSCENE (Yang et al. (2024b)) incorporates physical collision avoidance, object layouts, interactivity, and reachability metrics into a diffusion model. This approach enhances the physical plausibility and interactivity of generated scenes.
@@
Due to the prior knowledge of foundation models, there are current efforts to use foundation models for scene construction. RoboGen (Wang et al. (2023e)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023c)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library. However, using foundation models to automate the generation of scene’s physical plausibility still relies on VLM for judgment.
@@
The Real-to-Sim method and the Automated generation method both rely on 3D assets. The diversity of 3D assets determines the variety of scenes. RoboCasa (Nasiriany et al. (2024)) uses Madjourney to generate texture images for kitchen scenes. At the same time, many studies focus on the automatic generation of assets, such as zero-1-to-3 (Wang et al. (2023e)), Luma.ai (Nasiriany et al. (2024)), LLaMA-Mesh (Wang et al. (2024d)), Trellis (Xiang et al. (2024)). The modeling of the interaction environment above primarily focuses on articulated object modeling. Articulated objects can be created manually by designers or generated using procedural (Jiang et al. (2022); Liu et al. (2023b); Zhang et al. (2023e)) or human-assisted interactive methods (Torne et al. (2024)) after 3D scanning. They can also be generated automatically through inverse graphics (Chen et al. (2024c)). However, current automated methods for generating articulated object assets are limited to objects with few rotational joints. Real2Code (Mandi et al. (2024)) fine-tunes a CodeLlama model to process visual observation descriptions and then outputs joint predictions. This enables Real2Code to reconstruct complex articulated objects with up to 10 parts.

### v5

@@ 7.4 Foundation Models assisting for Reinforcement Learning
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. ROBOFUME (Yang et al. (2024a)) employs a fine-tuned VLM as the sparse reward function for the RL algorithm, tackling the issue of the extensive human supervision needed for training or fine-tuning a policy in the real world. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function. FAC (Ye et al. (2023a)) proposes using knowledge from foundation models as policy prior knowledge to improve sampling action efficiency, as value prior knowledge to measure the values of states and as success-reward prior knowledge to provide final feedback on task success.
@@
According to Fig. 7, policies can be classified into VLAC, VLAKP, VLADP, and Foundation Models assisting for Reinforcement Learning. As shown in Appendix.A Tab. 4 and Tab. 5, comparison with baseline approaches reveals key distinctions. The strengths and limitations of VLAC, VLAKP, and VLADP are as described in Tab. 1.
@@ 8 Manipulation Data Generation
Demonstration data plays a crucial role in robotic manipulation, particularly in the context of imitation learning (Padalkar et al. (2023b)). A common approach for gathering such demonstrations is human teleoperation in the real world. However, collecting real-world data often requires significant human labor and specialized teleoperation equipment. Recently, there has been a growing number of excellent developments in low-cost teleoperation hardware, which enables the collection of high-quality demonstration data (Fang et al. (2024); Cheng et al. (2024)).
@@
To collect data in real environments, human effort is required for scene setup and data annotation (Sermanet et al. (2023)). There are currently two methods for data collection: the bottom-up approach and the top-down step-by-step approach. The bottom-up approach focuses on selecting a task to perform based on the current scene. Then, it uses methods like crowd-sourcing to label the data. The top-down approach involves a step-by-step process where decision-makers assign task labels and manage overheads, such as resets and scene preparations (Sermanet et al. (2023)). The robot then performs tasks according to these labels. RoboVQA (Sermanet et al. (2023)) shows that the bottom-up approach is more efficient in data collection compared to the top-down step-by-step approach. DIAL (Xiao et al. (2022a)) uses a fine-tuned CLIP to replace humans in labeling robot trajectories during bottom-up data collection. This transforms the robot manipulation dataset on the internet into the robot-language manipulation dataset. PAFF (Ge et al. (2023)) points out that incorrect robot trajectories can be linked to new tasks and uses fine-tuned CLIP to label the incorrect robot trajectories with appropriate task labels. The above methods demonstrate that high-level cognitive models can assist in data annotation. SOAR (Zhou et al. (2024b)) shows that integrating a high-level cognitive model with a low-level control policy can result in a fully autonomous data-collection system in varied real-world environments.
@@
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023c)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge.
Compared to single-frame images and language data on the internet, internet videos contain information on the physics and dynamics of the world, as well as on human behaviors and actions (Chandrasegaran et al. (2024)). This information is precisely what is required for manipulation tasks. Therefore, in this section, we also introduce the internet-scale video data for robot learning.
Regardless of whether it’s in a real or simulated environment, improving the efficiency of the existing dataset is essential. The mainstream approach is dataset augmentation.
Table 2: Representative Low-cost Hardware Works.
Teleoperation Device Teleoperation Type Human Motion Measurement Feedback Cost Embodiment Config Manipulation Task
Aloha(Zhao et al. (2023a)) Online Teleoperation Joint Copy Third-person View \$2000(include robot) Dual-Arm; 2F Gripper
Slide Ziploc;
Slot Battery;
Open Cup; etc;
GELLO(Wu et al. (2023c)) Online Teleoperation Joint copy Third-person View \$300 Dual-Arm; 2F Gripper
Place Hat on Rack;
Hand Over;
Fold Towel; etc;
Human Plus(Fu et al. (2024)) Online Teleoperation End-effectors mapping Third-person View \$30(A RGB Camera) Humanoid Robot; Dexterous Hand
Wear Shoe and Walk;
Fold Clothes;
Warehouse; etc;
Transteleop(Li et al. (2020)) Online Teleoperation End-effectors mapping Third-person View \$170 Dual-Arm; Dexteours Hand
Pick-Place;
Handover; Push; etc;
OPEN TEACH(Iyer et al. (2024)) Online Teleoperation End-effectors mapping First-person View \$500 Dual-Arm; Dexteours Hand/2F Gripper;
Make Sandwich;
Ironing Cloth; Open Cabinet; etc;
Open-TeleVision(Cheng et al. (2024)) Online Teleoperation End-effectors mapping First-person View \$3499 Humanoid Robot; Dexteours Hand
Can Sorting;
Can Insertion;
Folding;
Unloading
AirExo (Fang et al. (2024)) Offline Teleoperation Joint copy - \$600 Dual-Arm; 2F Gripper
Gather Balls;
Grasp from the Curtained Shelf;
UMI(Chi et al. (2024)) Offline Teleoperation End-effectors mapping - \$371 Dual-Arm; 2F Gripper
Dish Washing;
Dynamic Tossing; Cloth Folding; etc;
DexCap(Wang et al. (2024b)) Offline Teleoperation End-effectors mapping - \$4000 Dual-Arm; Dexterous Hand
Scissor Cutting;
Tea Preparing; Sponge Picking; etc;
VideoDex(Shaw et al. (2023)) Offline Teleoperation End-effectors mapping - - Single-Arm; Dexterous Hand
Pick-Place;
Cover and Uncover;etc;
8.1 Low-cost Teleoperation Device
Above systems work in cartesian space, which needs inverse kinematic (IK) solver and off-the-shelf IK often suffering from fails when operating near singularities of the robot. Although some bilateral teleoperation systems use haptic feedback to provide a tangible sense of the robot’s kinematic constraints, they do not address the challenges of very tight operational spaces (Silva et al. (2009)). Therefore, multi-joint encoder teleoperation devices can solve the IK problem by working in the joint space. The current design of multi-joint teleoperation devices is mainly divided into isomorphic and non-isomorphic devices (Wu et al. (2023c)). Isomorphic devices refer to teleoperation systems using standard servo-based robotic arms to control manipulators with similar size and kinematics (Zhao et al. (2023a)), while non-isomorphic devices use such arms to control manipulators with different size and kinematic properties. Non-isomorphic devices use kinematically equivalent structures based on DH parameters to map joint spaces between different properties (Wu et al. (2023c)). AirExo (Fang et al. (2024)) expands this low-cost and scalable platform into a wearable device to collect cheap in-the-wild demonstrations at scale.
@@ 8.2 Simulator
The current mainstream simulators (Zhou et al. (2023)) include PyBullet (Coumans and Bai (2016)), MuJoCo (Todorov et al. (2012)), CoppeliaSim (Rohmer et al. (2013)), NVIDIA Omniverse and Unity. Pybullet is easy to use and integrate, but its graphics are quite basic. It is not suitable for applications that require complex visual effects. Therefore, Pybullet is often used together with Blender (Shi et al. (2024)). Mujoco offers a high-precision physics engine. It is suitable for simulating articulated and deformable object manipulation. However, it has a high entry barrier for beginners. CoppeliaSim offers a wide range of ready-made environments, objects, and prototyping robotic systems for users. However, when dealing with many robots or complex scenes, CoppeliaSim may encounter performance issues. NVIDIA Omniverse provides real-time physics simulation and realistic rendering. However, it requires significant computational resources. NVIDIA Omniverse offers many interfaces. Users can use these to develop various applications. For example, Issac Gym is a platform for robot reinforcement learning, developed using Omniverse. Unity offers rich visual effects and a user-friendly interface. It allows for the creation of highly interactive applications. However, its physics engine is still not precise enough. The basic components of a simulator are the physics engine and the renderer. Improvements in these components can enhance the capability of sensors in simulations, such as optical tactile sensors (Chen et al. (2023d)). Learning-based simulators also show great potential. For example, Sora (Brooks et al. (2024)) and UniSim (Yang et al. (2023b)) use vast amounts of data from the internet to simulate the visual effects of many different actions.
@@ 8.3 Scene and Demonstration Generation
The Real-to-Sim method directly refers to a digital twin. The Real-to-Sim method utilizes 3D-reconstruction technology (in Sec. 6) or inverse graphics (Chen et al. (2024c)) to create the real-world scene in a virtual environment (Torne et al. (2024)). But, 3D reconstruct scene is static environment where objects lack real-world physical properties, such as material, mass and friction coefficients, and are non-interactive (Torne et al. (2024)). The inverse graphics method, such as URDFormer (Chen et al. (2024c)), directly generates interactive simulation environment and articulated objects from input RGB image. Compared to 3D-reconstruction methods, it reduces human involvement and produces interactive simulation environment. However, it lacks physical plausibility and fails to address the mismatch between the generated object’s physical properties and the real world.
@@
As for the application of foundation models in Real-to-Sim methods, GRS (Zook et al. (2024)) employs SAM2 for object segmentation from RGBD image and utilizes VLMs to describe and match objects with simulation-ready assets. This approach combines the strengths of 3D-reconstruction and inverse graphics methods. It ensures the credibility of 3D-reconstruction methods and allowing objects in the scene to interact. However, it is impossible for the assert dataset to fully cover objects in the real world. Constructing an interact assert dataset often requires manual design by the creator or human-assisted interactive object generation. ACDC (Dai et al. (2024)) defines a digital cousin concept. Unlike a digital twin, it does not directly replicate a real-world counterpart. However, it retains similar geometric and semantic features by using similar asserts when the assert dataset does not include real-world objects. As for object pose, depth cameras are commonly used, but they struggle to capture reflective surfaces accurately. This limits their use in the wild. To address this issue, ACDC uses Depth-Anything-v2 (Yang et al. (2024b)), a state-of-the-art monocular depth estimation model, to estimate the depth map.
@@
Scene diversity primarily includes the diversity of scene layouts, such as floor plans and object placements, as well as the diversity of objects. The automated generation methods are more effective for producing large-scale diverse scenes. The automated generation methods can be categorized into rule-based and learning-based approaches. For instance, ProcTHOR (Deitke et al. (2022)) introduces a procedural generation pipeline for interactive scenes using rule-based constraints and statistical priors. However, the generated scenes often rely on pre-defined priors, resulting in unrealistic outcomes that hinder agent learning (Khanna et al. (2024)). To address this, PHYSCENE (Yang et al. (2024c)) incorporates physical collision avoidance, object layouts, interactivity, and reachability metrics into a diffusion model. This approach enhances the physical plausibility and interactivity of generated scenes.
@@
Due to the prior knowledge of foundation models, there are current efforts to use foundation models for scene construction. RoboGen (Wang et al. (2023e)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023c)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library. However, using foundation models to automate the generation of scene’s physical plausibility still relies on VLM for judgment. At the same time, the above research also uses LLMs to generate diverse instructions to ensure task diversity. However, generating diverse task instructions with LLMs presents challenges in ensuring rationality for the current environment.
The Real-to-Sim method and the Automated generation method both rely on 3D assets. The diversity of 3D assets determines the variety of scenes (Nasiriany et al. (2024)). Although there are many existing 3D object assets (Chang et al. (2015); Deitke et al. (2023); Li et al. (2023a); Geng et al. (2023a); Xiang et al. (2020); Liu et al. (2022b); Calli et al. (2017)), their quantity is far from sufficient to cover the variety of real-world objects. As a result, many studies focus on the automatic generation of assets, such as zero-1-to-3 (Wang et al. (2023e)), Luma.ai (Nasiriany et al. (2024)), LLaMA-Mesh (Wang et al. (2024d)), Trellis (Xiang et al. (2024)). However, the performance of generative models is also limited by the shortage of current 3D training data. To address this issue, data cleaning techniques or manual supervision are needed to filter and select high-quality generated object assets.
@@
The modeling of the interaction environment above primarily focuses on articulated object modeling. Articulated objects can be created manually by designers or generated using procedural (Jiang et al. (2022); Liu et al. (2023b); Zhang et al. (2023e)) or human-assisted interactive methods (Torne et al. (2024)) after 3D scanning. They can also be generated automatically through inverse graphics (Chen et al. (2024c)) or generative model (Xiang et al. (2024)). However, current automated methods for generating articulated object assets are limited to objects with few rotational joints. Real2Code (Mandi et al. (2024)) fine-tunes a CodeLlama model to process visual observation descriptions and then outputs joint predictions. This enables Real2Code to reconstruct complex articulated objects with up to 10 parts. At the same time, generative models mainly focus on rigid and articulated objects and research on deformable objects remains insufficient (Sundaresan et al. (2022)).